# Multi-orbital two-particle self-consistent approach – strengths and limitations

### v1

@@ Abstract
Extending many-body numerical techniques which are powerful in the context of simple model calculations to the realm of realistic material simulations can be a challenging task. Realistic systems often involve multiple active orbitals, which increases the complexity and numerical cost because of the large local Hilbert space and the large number of interaction terms or sign-changing off-diagonal Green’s functions. The two-particle self-consistent approach (TPSC) is one such many-body numerical technique, for which multi-orbital extensions have proven to be involved due to the substantially more complex structure of the local interaction tensor. In this paper we extend earlier multi-orbital generalizations of TPSC by setting up two different variants of a fully self-consistent theory for TPSC in multi-orbital systems. We first investigate the strengths and limitations of the approach analytically, and then benchmark both variants against exact diagonalization (ED) and dynamical mean field theory (DMFT) results. We find that the exact behavior of the system can be faithfully reproduced in the weak coupling regime, while at stronger couplings the performance of the two approaches strongly depends on details of the system.
preprint: APS/123-QED
I Introduction
Solving the many-body problem of interacting electrons remains a central challenge in condensed matter physics. Analytical solutions can only be found in rare cases [1, 2], necessitating the development of advanced numerical methods such as Density Functional Theory [3, 4], Quantum Monte Carlo (QMC) [5, 6, 7, 8, 9, 10], DMRG and other Tensor network approaches [11, 12, 13], Dynamical Mean-Field Theory and its extensions [14, 15, 16, 17, 18, 19], variational (neural) quantum states [20, 21, 22, 10, 23, 24, 25] or various diagrammatic methods [26, 27, 28, 29, 30, 31]. With the use of these approaches the community gained insights into correlated electron physics, e.g. by developing an understanding of Mott physics [32, 33, 34, 35], (unconventional) superconductivity [36, 37, 38, 39, 40], the pseudogap phase [41, 42], magnetism and spin liquids [43, 44, 45] and charge ordered states [46, 47, 48]. In recent years, a coherent picture has emerged for some single-orbital models, with a variety of numerical approaches producing consistent results for ground state energies, mass renormalizations, etc. [49, 50, 51], which shifts the frontiers of method development to more complex models and setups.
@@ I Introduction
Numerical studies can provide guiding principles and insights into the microscopic mechanisms behind emergent phenomena. However, in cases where the employed numerical approach fails qualitatively or quantitatively, as for example in the case of some iron-based superconductors where neither DFT nor DFT+DMFT predict correct Fermi surfaces [65, 66, 67], it is important to extend the techniques and improve their accuracy. A promising route for the study of correlated materials is to extend the formalism beyond DMFT, which is a dynamical but local approach, by combining it with another numerical approach which captures the effects of spatial and temporal fluctuations [14]. For this, two different strategies have been considered in recent years. One idea is to include finite distance correlations in the DMFT itself, at the cost of a more complicated impurity problem, as in the case of cellular DMFT [68, 69, 70] and the dynamical cluster approximation [71, 72]. Alternatively, one can extend the locality-approximation of DMFT from the self-energy to some vertex-functions, as done for example in (D-)TRILEX [18, 19], the dynamical-vertex approximation [17, 73] and other schemes [14, 74, 75, 76]. These approaches however typically require the calculation of local vertex functions, which is a time consuming and challenging task for complicated models [77].
@@
The paper is structured as follows: In Section II, we derive the central equations of TPSC from scratch and analytically discuss potential shortcomings. In Section III.1, we analyze a density-density interaction-only model analytically within TPSC, and point out potential pitfalls. In Section III.2 we analyze the effects of a strong Hunds-coupling. Next, in Section III.3, we benchmark the accuracy of the self-consistently determined two-particle expectation values, by comparing the TPSC results to exact diagonalization (ED) and DMFT calculations. In addition, we compare the spin and charge susceptibilities to D-TRILEX [103] and discuss the implications of these results for the applicability of our TPSC formulations to realistic multi-orbitals models.
II Multi-orbital TPSC
@@
$H=t_{o_{1}^{\prime},o_{3}^{\prime}}(\bm{r}_{1}^{\prime}-\bm{r}_{3}^{\prime})c^{\dagger}_{o_{3}^{\prime},s^{\prime}}(\bm{r}^{\prime}_{3})c_{o_{1}^{\prime},s^{\prime}}(\bm{r}_{1}^{\prime})+\frac{1}{2}U^{s_{1}^{\prime}s_{2}^{\prime}s_{3}^{\prime}s_{4}^{\prime}}_{o_{1}^{\prime}o_{2}^{\prime}o_{3}^{\prime}o_{4}^{\prime}}c^{\dagger}_{o_{3}^{\prime},s_{3}^{\prime}}(\bm{r}^{\prime})c^{\dagger}_{o_{4}^{\prime},s_{4}^{\prime}}(\bm{r}^{\prime})c_{o_{2}^{\prime},s_{2}^{\prime}}(\bm{r}^{\prime})c_{o_{1}^{\prime},s_{1}^{\prime}}(\bm{r}^{\prime}),~{}$ (1)
where $c^{(\dagger)}_{o_{1},s_{1}}(\tau_{1},\bm{r}_{1})$ annihilates (creates) an electron at orbital $o_{1}$ with spin $s_{1}$ at the lattice position $\bm{r}_{1}$. All primed variables are summed over. As a short hand for all non explicitly stated quantum numbers of an object we will use plain numbers as indices. We will restrict ourselves to the case of an $SU(2)$ symmetric model. To reduce the complexity, we rewrite the Hubbard interaction tensor in terms of its even and odd $SU(2)$-transforming components [104]
$U^{s_{1}s_{2}s_{3}s_{4}}_{o_{1}o_{2}o_{3}o_{4}}=U_{o_{1}o_{2}o_{3}o_{4}}\delta_{s_{1},s_{3}}\delta_{s_{2},s_{4}}-U_{o_{1}o_{2}o_{4}o_{3}}\delta_{s_{1},s_{4}}\delta_{s_{2},s_{3}}\;.$ (2)
Lastly, we restrict ourselves to inter-orbital-bilinear type interactions [105] (still allowing for intra-orbital interactions within this bilinear form), thus the spin independent interaction tensor is simplified to
@@
where each of the contributions can be identified with a specific physical process: density-density type interactions are given by $D$, spin-flip interactions by $C$ and pair-hopping processes by $P$. TPSC can also be formulated without this restriction, however restricting the interaction allows for an easier understanding of the equations and processes involved. Thus we focus on this simplified form.
@@
The central equation on which TPSC is built upon is the equation of motion [78], linking the product of the single particle self-energy and the Green’s function to the tensor contraction of the two-particle interaction with a two-particle expectation value (TPEV), which can be re-expressed in terms of a generalized susceptibility. For a derivation of the equation of motion, see Appendix A. With the assumption of an $SU(2)$-symmetric inter-orbital bilinear-type interaction, the equation of motion simplifies to
$\Sigma_{1,1^{\prime}}G_{1^{\prime},5}$ $=\frac{1}{2}U_{o_{2}^{\prime}o_{3}^{\prime}o_{4}^{\prime}o_{1}}^{s_{2}^{\prime}s_{3}^{\prime}s_{4}^{\prime}s_{1}}\braket{\mathcal{T}c^{\dagger}_{4^{\prime}}(\tau_{1},r_{1})c_{3^{\prime}}(\tau_{1},r_{1})c_{2^{\prime}}(\tau_{1},r_{1})c_{5}^{\dagger}}$ (4)
$=\frac{1}{2}U_{o_{2}^{\prime},o_{3}^{\prime},o_{4}^{\prime},o_{1}}\braket{\mathcal{T}c_{o_{4}^{\prime},s_{4}^{\prime}}^{\dagger}(\tau_{1},r_{1})c_{o_{3}^{\prime},s_{1}}(\tau_{1},r_{1})c_{o_{2}^{\prime},s_{4}^{\prime}}(\tau_{1},r_{1})c_{5}^{\dagger}}$ (5)
$\quad-\frac{1}{2}U_{o_{2}^{\prime},o_{3}^{\prime},o_{1},o_{4}^{\prime}}\braket{\mathcal{T}c_{o_{4}^{\prime},s_{4}^{\prime}}^{\dagger}(\tau_{1},r_{1})c_{o_{3}^{\prime},s_{4}^{\prime}}(\tau_{1},r_{1})c_{o_{2}^{\prime},s_{1}}(\tau_{1},r_{1})c_{5}^{\dagger}}$
$=U_{o_{2}^{\prime},o_{3}^{\prime},o_{4}^{\prime},o_{1}}\braket{\mathcal{T}c_{o_{4}^{\prime},s_{4}^{\prime}}^{\dagger}c_{o_{3}^{\prime},s_{1}}c_{o_{2}^{\prime},s_{4}^{\prime}}c_{5}^{\dagger}}\bar{\delta}_{s_{1},s_{4}^{\prime}}\bar{\delta}_{o_{2}^{\prime},o_{3}^{\prime}}$ (6)
$=\Big{(}D_{o_{4}^{\prime},o_{1}}\sum_{s_{4}^{\prime}}\bar{\delta}_{s_{1},s_{4}^{\prime}}\bar{\delta}_{o_{1},o_{4}^{\prime}}\braket{\mathcal{T}c_{o_{4}^{\prime},s_{4}^{\prime}}^{\dagger}c_{o_{1},s_{1}}c_{o_{4}^{\prime},s_{4}^{\prime}}c_{5}^{\dagger}}+C_{o_{1},o_{4}^{\prime}}\sum_{s_{4}^{\prime}}\bar{\delta}_{s_{1},s_{4}^{\prime}}\bar{\delta}_{o_{1},o_{4}^{\prime}}\braket{\mathcal{T}c_{o_{4}^{\prime},s_{4}^{\prime}}^{\dagger}c_{o_{4}^{\prime},s_{1}}c_{o_{1},s_{4}^{\prime}}c_{5}^{\dagger}}$ (7)
$\quad+P_{o_{2}^{\prime},o_{1}}\sum_{s_{4}^{\prime}}\bar{\delta}_{s_{1},s_{4}^{\prime}}\braket{\mathcal{T}c_{o_{1},s_{4}^{\prime}}^{\dagger}c_{o_{2}^{\prime},s_{1}}c_{o_{2}^{\prime},s_{4}^{\prime}}c_{5}^{\dagger}}\Big{)}\,,$
where we introduced $\bar{\delta}$ as a short hand for $1-\delta$. The factor of $2$ cancelling the $1/2$ stems from applying the remaining crossing symmetry (exchanging in-going and out-going indices at the same time), then swapping two operators and renaming summation indices. The operator swap results in the $1-\delta$. The equation of motion relates single particle to two particle properties, which in turn are linked to three particle properties. Therefore, the set of equations is not amenable to an exact solution due to this hierarchical structure of the equations.
@@
Instead, we proceed by approximating the two-particle expectation values on the right-hand side by their Hartree-Fock decoupling. However, to improve over Hartree-Fock, in TPSC one introduces parameters for the prefactors which are chosen such that the local and static limit of the two-particle expectation values are exactly recovered [78]. The Hartree-Fock decoupling reads
$\Sigma_{o_{1},o_{1}^{\prime}}^{s_{1},s_{1}^{\prime}}(\tau_{1},r_{1};\tau_{1}^{\prime},r_{1}^{\prime})G_{o_{1}^{\prime},o_{5}}^{s_{1}^{\prime},s_{5}}(\tau_{1}^{\prime},r_{1}^{\prime};\tau_{5},r_{5})$ $\approx\Big{(}\tilde{D}_{o_{4}^{\prime},o_{1}}^{s_{4}^{\prime},s_{1}}(n_{o_{4}^{\prime}}^{s_{4}^{\prime}}G_{o_{1},o_{5}}^{s_{1},s_{5}}(\tau_{1},r_{1};\tau_{5},r_{5})-G_{o_{1},o_{4}^{\prime}}^{s_{1},s_{4}^{\prime}}G_{o_{4}^{\prime},o_{5}}^{s_{4}^{\prime},s_{5}}(\tau_{1},r_{1};\tau_{5},r_{5}))$
$\quad\tilde{C}^{s_{1},s_{4}^{\prime}}_{o_{1},o_{4}^{\prime}}(G_{o_{4}^{\prime},o_{4}^{\prime}}^{s_{1},s_{4}^{\prime}}G_{o_{1},o_{5}}^{s_{4}^{\prime},s_{5}}(\tau_{1},r_{1};\tau_{5},r_{5})-G_{o_{1},o_{4}^{\prime}}^{s_{4}^{\prime},s_{4}^{\prime}}G_{o_{4}^{\prime},o_{5}}^{s_{1},s_{5}}(\tau_{1},r_{1};\tau_{5},r_{5}))$
$\quad+\tilde{P}^{s_{1},\bar{s}_{1}}_{o_{2}^{\prime},o_{1}}(G_{o_{2}^{\prime},o_{1}}^{s_{1},\bar{s}_{1}}G_{o_{2}^{\prime},o_{5}}^{\bar{s}_{1},s_{5}}(\tau_{1},r_{1};\tau_{5},r_{5})-G_{o_{2}^{\prime},o_{1}}^{\bar{s}_{1},\bar{s}_{1}}G_{o_{2}^{\prime},o_{5}}^{s_{1},s_{5}}(\tau_{1},r_{1};\tau_{5},r_{5}))\Big{)}\,,$ (8)
where $\bar{s}=-s$.
It should be noted that there is an ambiguity in the way we define the prefactors: we can either define them as spin dependent or spin independent quantities – in the latter case, we end up with three independent vertices, while in the former case, we end up with five. In the following we will refer to these two variants as TPSC3 and TPSC5, respectively.

### v3

@@ Abstract
Extending many-body numerical techniques which are powerful in the context of simple model calculations to the realm of realistic material simulations can be a challenging task. Realistic systems often involve multiple active orbitals, which increases the complexity and numerical cost because of the large local Hilbert space and the large number of interaction terms or sign-changing off-diagonal Green’s functions. The two-particle self-consistent approach (TPSC) is one such many-body numerical technique, for which multi-orbital extensions have proven to be involved due to the substantially more complex structure of the local interaction tensor. In this paper we extend earlier multi-orbital generalizations of TPSC by setting up two different variants of a fully self-consistent theory for TPSC in multi-orbital systems. We first investigate the strengths and limitations of the approach analytically and then benchmark both variants against dynamical mean-field theory (DMFT) and D-TRILEX results. We find that the exact behavior of the system can be faithfully reproduced in the weak-coupling regime, while at stronger couplings the performance of the two TPSC variants strongly depends on details of the system.
Copyright attribution to authors.
This work is a submission to SciPost Physics Lecture Notes.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
1 Introduction
Solving the many-body problem of interacting electrons remains a central challenge in condensed matter physics. Analytical solutions can only be found in rare cases [1, 2], necessitating the development of advanced numerical methods such as Density Functional Theory [3, 4], Quantum Monte Carlo (QMC) [5, 6, 7, 8, 9, 10], DMRG and other Tensor network approaches [11, 12, 13], Dynamical Mean-Field Theory and its extensions [14, 15, 16, 17, 18, 19], variational (neural) quantum states [20, 21, 22, 10, 23, 24, 25] or various diagrammatic methods [26, 27, 28, 29, 30, 31]. With the use of these approaches the community gained insights into correlated electron physics, e.g. by developing an understanding of Mott physics [32, 33, 34, 35], (unconventional) superconductivity [36, 37, 38, 39, 40], the pseudogap phase [41, 42], magnetism, spin liquids [43, 44, 45] and charge ordered states [46, 47, 48]. In recent years, a coherent picture has emerged for some single-orbital models, with a variety of numerical approaches producing consistent results for ground state energies, mass renormalizations, etc. [49, 50, 51], which shifts the frontiers of method development to more complex models and setups.
@@ 1 Introduction
Numerical studies can provide guiding principles and insights into the microscopic mechanisms behind emergent phenomena. However, in cases where the employed numerical approach fails qualitatively or quantitatively, as for example in the case of some iron-based superconductors where neither DFT nor DFT+DMFT predict correct Fermi surfaces [65, 66, 67], it is important to extend the techniques and improve their accuracy. A promising route for the study of correlated materials is to extend the formalism beyond DMFT, which is a dynamical but local approach, by combining it with another numerical approach which captures the effects of spatial and temporal fluctuations [14]. For this, two different strategies have been considered in recent years. One idea is to include finite distance correlations in the DMFT itself, at the cost of a more complicated impurity problem, as in the case of cellular DMFT [68, 69, 70] and the dynamical cluster approximation [71, 72]. Alternatively, one can extend the locality-approximation of DMFT from the self-energy to some vertex-functions, as done for example in TRILEX [18] and D-TRILEX [73, 74], the dynamical-vertex approximation [17, 75] and other schemes [14, 76, 77, 78]. These approaches however typically require the calculation of local vertex functions, which is a time consuming and challenging task for complicated models [79].
@@
The paper is structured as follows: In Section 2, we derive the central equations of multi-orbital TPSC from scratch. In Section 3.1, we analyze a density-density interaction-only model analytically within TPSC, and point out potential pitfalls. In Section 3.2 we analyze the effects of a strong Hunds coupling. Next, in Section 3.3, we benchmark the accuracy of the self-consistently determined two-particle expectation values, by comparing the TPSC results to DMFT calculations. In addition, we compare the self-energy, spin and charge susceptibilities to D-TRILEX [19] and discuss the implications of these results for the applicability of our TPSC formulations to realistic multi-orbitals models.
2 Multi-orbital TPSC
@@
$H=t_{o_{1^{\prime}},o_{3^{\prime}}}(\boldsymbol{r}_{1}^{\prime}-\boldsymbol{r}_{3}^{\prime})$ $c^{\dagger}_{o_{3^{\prime}},s^{\prime}}(\boldsymbol{r}^{\prime}_{3})c_{o_{1^{\prime}},s^{\prime}}(\boldsymbol{r}_{1}^{\prime})$ (1)
$+\frac{1}{2}\tilde{V}^{s_{1^{\prime}}s_{2^{\prime}}s_{3^{\prime}}s_{4^{\prime}}}_{o_{1^{\prime}}o_{2^{\prime}}o_{3^{\prime}}o_{4^{\prime}}}c^{\dagger}_{o_{3^{\prime}},s_{3^{\prime}}}(\boldsymbol{r}^{\prime})c^{\dagger}_{o_{4^{\prime}},s_{4^{\prime}}}(\boldsymbol{r}^{\prime})c_{o_{2^{\prime}},s_{2^{\prime}}}(\boldsymbol{r}^{\prime})c_{o_{1^{\prime}},s_{1^{\prime}}}(\boldsymbol{r}^{\prime}),~{}$
where $c^{(\dagger)}_{o_{1},s_{1}}(\boldsymbol{r}_{1})$ annihilates (creates) an electron at orbital $o_{1}$ with spin $s_{1}$ at the lattice position $\boldsymbol{r}_{1}$. $t_{o_{1},o_{3}}(\boldsymbol{r}_{1}-\boldsymbol{r}_{3})$ is the hopping matrix element between two orbitals at distance $\boldsymbol{r}_{1}-\boldsymbol{r}_{3}$ and $\tilde{V}^{s_{1}s_{2}s_{3}s_{4}}_{o_{1}o_{2}o_{3}o_{4}}$ is the local interaction matrix element. All primed variables are summed over. For the rest of the paper we will work in the Baym-Kadanoff formalism, thus the operators $c_{o_{1},s_{1}}(\boldsymbol{r})$ acquire a dependence on the imaginairy time $c_{o_{1},s_{1}}(\boldsymbol{r},\tau)$ We introduce the four vector $\boldsymbol{\tau}\equiv(\boldsymbol{r},\tau)$ as a short hand to compactify our notation. Furthermore, we define the anti-symmetrized interaction tensor
$V^{s_{1}s_{2}s_{3}s_{4}}_{o_{1}o_{2}o_{3}o_{4}}=\frac{1}{2}\mathcal{P}(\tilde{V}^{s_{1}s_{2}s_{3}s_{4}}_{o_{1}o_{2}o_{3}o_{4}})\,,$ (2)
where $\mathcal{P}$ is the anti-symmetrization operator. From now on, we will work with the antisymmetrized interation. In App. B we perform the analogous derivations starting from a Hubbard-Kanamori type instead of a general anti-symmetrical interaction tensor, which leads to an equivalent set of self-consistency equations. We will restrict ourselves to the case of an $SU(2)$ symmetric model. To reduce the complexity, we rewrite the interaction operator in terms of its even and odd $SU(2)$-transforming components [105]
$\quad{V}^{s_{1}s_{2}s_{3}s_{4}}_{o_{1}o_{2}o_{3}o_{4}}$ $=U_{o_{1}o_{2}o_{3}o_{4}}\delta_{s_{1},s_{3}}\delta_{s_{2},s_{4}}-{U}_{o_{1}o_{2}o_{4}o_{3}}\delta_{s_{1},s_{4}}\delta_{s_{2},s_{3}}\;.$ (3)
Lastly, we restrict ourselves to inter-orbital-bilinear type interactions [106] (still allowing for intra-orbital interactions within this bilinear form), thus the spin independent interaction tensor simplifies to
@@
where each of the contributions is native to a different diagram type: $D$ has the form of a resummation in the direct particle-hole channel, $C$ has the form of a crossed particle-hole one and a particle-particle resummation leads to terms of the form of $P$. TPSC can also be formulated without this restriction, however restricting the interaction allows for an easier understanding of the equations and processes involved.
@@
The central equation on which TPSC is built is the equation of motion [80], linking the product of the single particle self-energy and the Green’s function to the tensor contraction of the two-particle interaction with a two-particle expectation value (TPEV). The TPEV can then be re-expressed in terms of a generalized susceptibility. For a derivation of the equation of motion, see Appendix A up to Eq. (79). With the assumption of an $SU(2)$-symmetric inter-orbital bilinear-type interaction, the equation of motion simplifies to
$\Sigma_{o_{1},o_{1^{\prime}}}^{s_{1},s_{1^{\prime}}}$ $(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{1}^{\prime})G_{o_{1^{\prime}},o_{3}}^{s_{1^{\prime}},s_{3}}(\boldsymbol{\tau}_{1}^{\prime},\boldsymbol{\tau}_{3})=$ (5)
$\quad\frac{1}{2}V_{o_{4^{\prime}},o_{2^{\prime}},o_{3^{\prime}},o_{1}}^{s_{4^{\prime}},s_{2^{\prime}},s_{3^{\prime}},s_{1}}\braket{T_{\tau}c^{\dagger}_{o_{3^{\prime}},s_{3^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{2^{\prime}}s_{2^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{4^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3},s_{3}}^{\dagger}(\boldsymbol{\tau}_{3})}$
$=\frac{1}{2}U_{o_{4^{\prime}},o_{2^{\prime}},o_{3^{\prime}},o_{1}}\braket{T_{\tau}c^{\dagger}_{o_{3^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{2^{\prime}}s_{1}}(\boldsymbol{\tau}_{1})c_{o_{4^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3},s_{3}}^{\dagger}(\boldsymbol{\tau}_{3})}$ (6)
$\quad-\frac{1}{2}U_{o_{4^{\prime}},o_{2^{\prime}},o_{1},o_{3^{\prime}}}\braket{T_{\tau}c^{\dagger}_{o_{3^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{2^{\prime}}s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{4^{\prime}},s_{1}}(\boldsymbol{\tau}_{1})c_{o_{3},s_{3}}^{\dagger}(\boldsymbol{\tau}_{3})}$
$=U_{o_{4^{\prime}},o_{2^{\prime}},o_{3^{\prime}},o_{1}}\braket{T_{\tau}c^{\dagger}_{o_{3^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{2^{\prime}}s_{1}}(\boldsymbol{\tau}_{1})c_{o_{4^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3},s_{3}}^{\dagger}(\boldsymbol{\tau}_{3})}$ (7)
$=\Big{(}D_{o_{4^{\prime}},o_{1}}\sum_{s_{4^{\prime}}}\braket{T_{\tau}c_{o_{4^{\prime}},s_{4^{\prime}}}^{\dagger}(\boldsymbol{\tau}_{1})c_{o_{1},s_{1}}(\boldsymbol{\tau}_{1})c_{o_{4^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3},s_{3}}^{\dagger}(\boldsymbol{\tau}_{3})}$ (8)
$\quad+C_{o_{1},o_{3^{\prime}}}\sum_{s_{4^{\prime}}}\braket{T_{\tau}c^{\dagger}_{o_{3^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3^{\prime}}s_{1}}(\boldsymbol{\tau}_{1})c_{o_{1},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3},s_{3}}^{\dagger}(\boldsymbol{\tau}_{3})}$
$\quad+P_{o_{2^{\prime}},o_{1}}\sum_{s_{4^{\prime}}}\braket{T_{\tau}c^{\dagger}_{o_{1},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{2^{\prime}}s_{1}}(\boldsymbol{\tau}_{1})c_{o_{2^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3},s_{3}}^{\dagger}(\boldsymbol{\tau}_{3})}\Big{)}\,.$
The factor of $2$ canceling the $1/2$ stems from applying the remaining crossing symmetry (exchanging in-going and out-going indices at the same time), then swapping two operators and renaming summation indices. Furthermore, in the $P$ channel the only contribution to the spin sum which is non-zero is $s_{1}\neq s_{4}^{\prime}$ due to the Pauli-principle.
@@
The equation of motion relates single-particle to two-particle properties, which in turn are linked to three-particle properties. Therefore, the set of equations is not amenable to an exact solution due to this hierarchical structure of the equations. Instead, we proceed by approximating the two-particle expectation values on the right-hand side by their Hartree-Fock decoupling. However, to improve over Hartree-Fock, in TPSC one introduces parameters for the prefactors which are chosen such that the local and static limit of the two-particle expectation values fulfill exact local susceptibility sumrules [80]. In the multi-orbital case this means that we consider the limit $\boldsymbol{\tau}_{1}=\boldsymbol{\tau}_{3}$, $o_{1}=o_{3}$ and $s_{1}=s_{3}$. Technically, the choice of the Ansätze is completely arbitrary, here we restrict ourselves to the cases in which the different interaction terms do not mix. The Hartree-Fock decoupling with the Ansätze introduced then reads
$\Sigma_{o_{1},o_{1^{\prime}}}^{s_{1},s_{1^{\prime}}}$ $(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{1}^{\prime})G_{o_{1^{\prime}},o_{3}}^{s_{1^{\prime}},s_{3}}(\boldsymbol{\tau}_{1}^{\prime};\boldsymbol{\tau}_{3})$ (9)
$\approx\Big{(}\tilde{D}_{o_{4^{\prime}},o_{1}}^{s_{4^{\prime}},s_{1}}(G_{o_{4^{\prime}},o_{4^{\prime}}}^{s_{4^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{1})G_{o_{1},o_{3}}^{s_{1},s_{3}}(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{3})-G_{o_{1},o_{4^{\prime}}}^{s_{1},s_{4^{\prime}}}(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{1})G_{o_{4^{\prime}},o_{3}}^{s_{4^{\prime}},s_{3}}(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{3}))$
$\quad+\tilde{C}^{s_{1},s_{3^{\prime}}}_{o_{1},o_{3^{\prime}}}(G_{o_{1},o_{3^{\prime}}}^{s_{4^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{1})G_{o_{3^{\prime}},o_{3}}^{s_{1},s_{3}}(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{3})-G_{o_{3^{\prime}},o_{3^{\prime}}}^{s_{1},s_{4^{\prime}}}(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{1})G_{o_{1},o_{3}}^{s_{4^{\prime}},s_{3}}(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{3}))$
$\quad+\tilde{P}^{s_{1},\bar{s}_{1}}_{o_{2^{\prime}},o_{1}}(G_{o_{2^{\prime}},o_{1}}^{\bar{s}_{1},\bar{s}_{1}}(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{1})G_{o_{2^{\prime}},o_{3}}^{s_{1},s_{3}}(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{3})-G_{o_{2^{\prime}},o_{1}}^{s_{1},\bar{s}_{1}}(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{1})G_{o_{2^{\prime}},o_{3}}^{\bar{s}_{1},s_{3}}(\boldsymbol{\tau}_{1};\boldsymbol{\tau}_{3}))\Big{)}\,,$
where $\bar{s}=-s$. One important thing to keep in mind is that TPSC will partially inherit shortcomings of this underlying Hartree-Fock decoupling (we will discuss this in more detail in Sec 3.1). On the other hand, from Eq. (8) we find in the local and static limit
$\Sigma_{o_{1},o_{1^{\prime}}}^{s_{1},s_{1^{\prime}}}$ $(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{1}^{\prime})G_{o_{1^{\prime}},o_{1}}^{s_{1^{\prime}},s_{1}}(\boldsymbol{\tau}_{1}^{\prime},\boldsymbol{\tau}_{1})=$ (10)
$\Big{(}D_{o_{4^{\prime}},o_{1}}\sum_{s_{4^{\prime}}}\braket{T_{\tau}c_{o_{4^{\prime}},s_{4^{\prime}}}^{\dagger}(\boldsymbol{\tau}_{1})c_{o_{4^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{1},s_{1}}^{\dagger}(\boldsymbol{\tau}_{1})c_{o_{1},s_{1}}(\boldsymbol{\tau}_{1})}$
$\quad-C_{o_{1},o_{3^{\prime}}}\sum_{s_{4^{\prime}}}\braket{T_{\tau}c^{\dagger}_{o_{3^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3^{\prime}}s_{1}}(\boldsymbol{\tau}_{1})c_{o_{1},s_{1}}^{\dagger}(\boldsymbol{\tau}_{1})c_{o_{1},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})}$
$\quad-P_{o_{2^{\prime}},o_{1}}\braket{T_{\tau}c^{\dagger}_{o_{1},\bar{s}_{1}}(\boldsymbol{\tau}_{1})c_{o_{2^{\prime}}s_{1}}(\boldsymbol{\tau}_{1})c_{o_{1},s_{1}}^{\dagger}(\boldsymbol{\tau}_{1})c_{o_{2^{\prime}},\bar{s}_{1}}(\boldsymbol{\tau}_{1})}\Big{)}\,.$
By combining Eq. (9) and Eq. (10) we obtain the explicit form of the Ansätze. It should be noted that there is an ambiguity in the way we define the prefactors: we can either define them as objects depending on one or on two spin indices. In the former case, we end up with three independent vertices, while in the latter case, we end up with five. The ansatz proportional to $P$ always depends only on a single spin. The origin of this ambiguity is the freedom of pulling the sum over the internal spin in Eq. (8) into the Ansatz (leading to TPSC3) or performing the Ansatz inside the sum (leading to TPSC5). Whether both of these choices are valid in the sense that no artificial $SU(2)$ symmetry breaking is introduced has to be checked by deriving the interaction vertex induced by them. Again we stress that there is technically a much larger group of possible Ansätze due to the freedom of shifting contributions to different interaction terms.